OpenStudy (anonymous):

perform the indicated operations & simplify 3z + 2z - 5 --- --- --- z^2-1 1-z z-1

OpenStudy (anonymous):

first you'll need to get the have the same denominator:\[\frac{3z + 2z(-(z+1)) - 5(z-1)(z+1)}{(z+1)} \]\[= \frac{3z -2z^2-2z-5z^2+5}{(z+1)(z-1)}\]\[= \frac{z-7z^2+5}{(z-1)(z+1)}\]\[=\frac{-7z^2 +z +5}{(z-1)(z+1)}\] correct me if I'm wrong please ^_^

OpenStudy (anonymous):

first step (z-1)(z+1) <-denominator, typo lol ^_^"

OpenStudy (anonymous):

hmm, there's something wrong?

OpenStudy (anonymous):

I have \[(z(-2z-4)-5)/((z+1)(z-1)\]

OpenStudy (anonymous):

I think it has to do with the 1-z thing

OpenStudy (anonymous):

I think your solution has a wrong sign on the numerator?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

lol yeah, the 2z(-(1+z)) right|? :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

it should be 2z(z+1) ._. my bad lol sorry :)

OpenStudy (anonymous):

sorry that answer is not correct

OpenStudy (anonymous):

?